# SGD with Clipping is Secretly Estimating the Median Gradient

### v1

@@ A.3 Rates for Biased SGD for Convex Lipschitz Functions
$(\forall\bm{w}^{\prime}\in\mathbb{R}^{d})\ell(\bm{w}^{\prime})-\ell(\bm{w})-\langle\bm{u},\bm{w}^{\prime}-\bm{w}\rangle\geq 0.$
@@
Let $\ell:\mathbb{R}^{d}\to\mathbb{R}$ be a convex $G$-Lipschitz function such that ${\rm{argmin}}~{}\ell\neq\emptyset$. Let $\eta>0$, and note $D:=\|\bm{w}_{0}-\bm{w}_{*}\|$ for some $\bm{w}_{*}\in{\rm{argmin}}~{}\ell$. Consider the iterates generated by a biased Stochastic Subgradient Descent method
@@
$\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]\leq 2D\sqrt{\frac{\sigma^{2}+(\delta+G)^{2}}{T}+\frac{\delta^{2}}{2}}\quad\text{(resp. $\frac{2D^{2}+\sigma^{2}+\delta^{2}+G^{2}}{\sqrt{T}}+(2D^{2}+1)\delta$ ).}$.
@@
Let $w_{*}\in{\rm{argmin}}~{}\ell$ and $T\geq 1$ be fixed. Develop the squares and use the definition of the algorithm to write
@@
Let $\bar{\bm{g}}_{t}$ be the projection of $\mathbb{E}\left[\bm{g}_{t}|\mathcal{F}_{t}\right]$ onto $\partial\ell(\bm{w}_{t})$, the subdifferential of $\ell$ at $\bm{w}_{t}$. This means that from our assumption on the bias of $\bm{g}_{t}$ we will have
@@
Now we introduce a parameter $\eta>0$, and we use Young’s inequality together with our assumption on the bias of $\bm{g}_{t}$ to write
$2\langle\bar{\bm{g}}_{t}-\mathbb{E}_{t}\left[\bm{g}_{t}\right],\bm{w}_{t}-\bm{w}_{*}\rangle\leq\frac{1}{\eta}\|\bar{\bm{g}}_{t}-\mathbb{E}_{t}\left[\bm{g}_{t}\right]\|^{2}+\eta\|\bm{w}_{t}-\bm{w}_{*}\|^{2}\leq\eta^{-1}\delta^{2}+\eta\|\bm{w}_{t}-\bm{w}_{*}\|^{2}.$
@@
$\mathbb{E}_{t}\left[\|\bm{w}_{t+1}-\bm{w}_{*}\|^{2}\right]\leq(1+\eta\eta)\|\bm{w}_{t}-\bm{w}_{*}\|^{2}-2\eta(\ell(\bm{w}_{t})-\inf\ell)+\eta^{2}\sigma^{2}+\eta^{2}(\delta+G)^{2}+\eta^{-1}\delta^{2},$
@@
$\mathbb{E}\left[\ell(\bm{w}_{t})-\inf\ell\right]\leq\frac{1+\eta\eta}{2\eta}\mathbb{E}\left[\|\bm{w}_{t}-\bm{w}_{*}\|^{2}\right]-\frac{1}{2\eta}\mathbb{E}\left[\|\bm{w}_{t+1}-\bm{w}_{*}\|^{2}\right]+\eta\frac{\sigma^{2}+(\delta+G)^{2}}{2}+\eta^{-1}\frac{\delta^{2}}{2}.$
We are now ready to use a weighted telescopic sum argument. First, define $\theta:=\frac{1}{1+\eta\eta}$ and $\rho_{t}:=\theta^{t}$. Second, multiply our inequality by $\rho_{t+1}$ and then sum for $t=0,\dots,T-1$. Observe that the term $(1+\eta\eta)\rho_{t}$ is equal to $\rho_{t}$ due to our definition, which means that we have a telescoping sum where the terms $\rho_{t}\mathbb{E}\left[\|\bm{w}_{t}-\bm{w}_{*}\|^{2}\right]$ will cancel each other. Third, divide by $\sum_{t=0}^{T-1}\rho_{t+1}$ so to obtain
$\frac{1}{\sum_{t=0}^{T-1}\rho_{t+1}}\sum_{t=0}^{T-1}\rho_{t+1}\mathbb{E}\left[\ell(\bm{w}_{t})-\inf\ell\right]\leq\frac{\rho_{0}}{2\eta\sum_{t=0}^{T-1}\rho_{t+1}}\mathbb{E}\left[\|\bm{w}_{0}-\bm{w}_{*}\|^{2}\right]+\eta\frac{\sigma^{2}+(\delta+G)^{2}}{2}+\eta^{-1}\frac{\delta^{2}}{2}.$
@@
$\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]\leq\frac{\|\bm{w}_{0}-\bm{w}_{*}\|^{2}}{2\eta\sum_{t=0}^{T-1}\rho_{t+1}}+\eta\frac{\sigma^{2}+(\delta+G)^{2}}{2}+\eta^{-1}\frac{\delta^{2}}{2}.$
We will now simplify this upper bound, so we can later make an appropriate choice of $\alpha$ and $\eta$ leading to the desired complexity bound. We will focus now on the geometric sum, and start by using the definition of $\rho_{t}$ to write
@@
$\eta\frac{\theta}{1-\theta}=\frac{\eta}{1+\eta\eta}\frac{1}{1-\frac{1}{1+\eta\eta}}=\frac{\eta}{\eta\eta}=\eta^{-1}.$
On the other hand, we can guarantee that $(1-\theta^{T})\geq 1/2$ provided that $\eta=\frac{2}{\eta T}$. To see this, first observe that $(1-\theta^{T})\geq 1/2$ is equivalent to $T\geq\frac{\ln(2)}{\ln(\theta^{-1})}$. Since we impose that $\eta\eta=\frac{2}{T}$ then necessarily $\theta=\frac{T}{T+2}\Leftrightarrow\theta^{-1}=1+\frac{2}{T}$. So what we need is $T\geq\frac{\ln(2)}{\ln(1+\frac{2}{T})}$. Now, observe that $\ln(2)\leq 1$. Moreover, it is an exercise (see Lemma A.5 which is deferred to the end of this proof) to verify that for all $x\geq 0$, $\frac{1}{\ln(x+1)}\leq\frac{1}{2}+\frac{1}{x}$. So for our condition to hold it is enough to ask that $T\geq\frac{1}{2}+\frac{T}{2}$, which is equivalent to $T\geq 1$, which is true.
Combining all those bounds together with our new definition for $\eta$, we now have
@@
Minimizing the right-hand side is equivalent to solve $\frac{a}{\eta T}=\eta b+c\eta T$, where
@@
We see that $\phi^{\prime\prime}(x)\geq 0$ for all $x\geq 0$, so $\phi$ is convex on $[0,+\infty)$. So we can use the tangent inequality:
@@
A.4 Rates for SMGD for Convex Lipschitz Functions: Proof of Proposition 3.4
See 3.4
@@
We want to apply Theorem A.4. Let us denote $\bm{g}_{t}=\bm{m}_{t}+\bm{e}_{t}$, so that the iterates of SMGD verify $\bm{w}_{t+1}=\bm{w}_{t}-\eta\bm{g}_{t}$. We are now going to show that $\bm{g}_{t}$ has uniformly bounded bias and variance. For this we will make use of Assumptions 3.1 and 3.2. Remember that we assumed $\delta_{2}=\sigma_{2}=0$ without loss of generality (we can do this according to Lemma A.1), so we note $\delta$ and $\sigma$ instead of $\delta_{1}$ and $\sigma_{1}$. Also, Assumption 3.2 on the errors together with the tower rule imply that $\mathbb{E}_{t}\left[\|\bm{e}_{t}\|^{2}\right]=\mathbb{E}_{t}\left[\mathbb{E}\left[\|\bm{e}_{t}\|^{2}\ |\ \bm{g}_{t}^{(1)},\dots,\bm{g}_{t}^{(n)}\right]\right]\leq\varepsilon^{2}$. Regarding the bias, we can write
@@
Using Theorem A.4, we obtain
@@
A.5 Rates for SMGD with Stable Noise: Proof of Corollary 3.6
See 3.6
@@
First of all, remember that the $\ell_{1}$-median can be computed exactly, by computing component-wise a one-dimensional median. This allows us to have $\varepsilon=0$ in Assumption 3.2.
Proof of part (i). The result is a direct consequence of Lemma A.3 and Proposition 3.3.
Proof of part (ii). The result is a direct consequence of Lemma A.3 and Proposition 3.4.
@@
Appendix B Approximate Computation of Median-like Estimators
B.1 Median-like Estimators : General Considerations
@@
The function $\mathcal{D}:\mathbb{R}^{d}\to\mathbb{R}$ is convex, ${\rm{argmin}}~{}\mathcal{D}=\{0\}$ and the problem (4) admits a solution.
Note that assuming $\mathcal{D}$ to be convex and finite implies that it is continuous, therefore measurable, which means that $\mathbb{E}\left[\mathcal{D}(\bm{m}-\bm{z})\right]$ is well defined. The assumption that ${\rm{argmin}}~{}\mathcal{D}=\{0\}$ guarantees that in the trivial case when the distribution is a Dirac $\delta_{\bm{m}_{0}}$, then the solution to (4) is exactly $\bm{m}_{0}$. The assumption that a solution to (4) exists is easily verified in practice, under the assumption that $\mathcal{D}$ is coercive:
@@
Let $\mathcal{D}:\mathbb{R}^{d}\to\mathbb{R}$ be convex. Assume that $\mathbb{E}\left[\mathcal{D}(\bm{z}-\mathbb{E}\left[\bm{z}\right])\right]<+\infty$. If $\mathcal{D}$ is coercive, i.e. $\lim\limits_{\|m\|\to\infty}\mathcal{D}(\bm{m})=+\infty$, then (4) admits a solution.
@@
Let $\phi(\bm{m}):=\mathbb{E}\left[\mathcal{D}(\bm{m}-\bm{z})\right]$. Since we assume that $\phi(\mathbb{E}\left[\bm{z}\right])=\mathbb{E}\left[\mathcal{D}(\bm{z}-\mathbb{E}\left[\bm{z}\right])\right]<+\infty$ then $\phi$ is proper. Since $\mathcal{D}$ is convex then $\phi$ is convex, as an expectation of convex functions. Moreover $\mathcal{D}$ is convex and finite, which means that $\mathcal{D}$ is continuous on $\mathbb{R}^{d}$, see e.g. Corollary 8.39 in (Bauschke & Combettes, 2017). Therefore $\phi$ is an expectation of lower semicontinuous functions, which implies that $\phi$ is lower semicontinuous.
Since $\mathcal{D}$ is coercive, convex and continuous, we know that there exists $a\in(0,+\infty)$ and $b\in\mathbb{R}$ (see Proposition 14.16 in (Bauschke & Combettes, 2017)) such that $\mathcal{D}(\cdot)\geq a\|\cdot\|+b$. Therefore
@@
[…]
B.2 Approximations based on Subgradient Updates
Here, $\tau>0$ is the learning rate, and $\bm{u}_{t}$ is a (convex) subgradient. Since $\mathcal{D}$ could be a non-differentiable function such as the $\ell_{2}$- or $\ell_{1}$-norm, we use subgradients instead of gradients.
Squared $\ell_{2}$-norm.
[…]
$\ell_{1}$-norm.
If we choose $\mathcal{D}=\|\cdot\|_{1}$, update (26) gives
$\bm{m}_{t+1}=\bm{m}_{t}+\tau\operatorname{sgn}(\bm{g}_{t}-\bm{m}_{t}),$
where the $\operatorname{sgn}$-operator can take any value in $[-1,1]$ for coordinates $i\in[d]$ such that $(\bm{g}_{t}-\bm{m}_{t})_{i}=0$. This recovers a version of sign-SGD (Riedmiller & Braun, 1993; Bernstein et al., 2018), where the $\mathrm{sign}$-operation is applied to the increment $\bm{m}_{t}-\bm{g}_{t}$. For example, if we reset $\bm{m}_{t}=0$ in every iteration, we obtain exactly sign-SGD.
$\ell_{2}$-norm.
B.3 Approximations based on the Stochastic Proximal Point: General $\mathcal{D}$

### v2

@@ A.3 Rates for Biased SGD for Convex Lipschitz Functions
$(\forall\bm{w}^{\prime}\in\mathbb{R}^{d})\quad\ell(\bm{w}^{\prime})-\ell(\bm{w})-\langle\bm{u},\bm{w}^{\prime}-\bm{w}\rangle\geq 0.$
@@
Let $\ell:\mathbb{R}^{d}\to\mathbb{R}$ be a convex $G$-Lipschitz function such that $\rm{argmin}$$\ell\neq\emptyset$. Let $\eta>0$, and note $D:=\|\bm{w}_{0}-\bm{w}_{*}\|$ for some $\bm{w}_{*}\in{\rm{argmin}}$$\ell$. Consider the iterates generated by a biased Stochastic Subgradient Descent method
@@
$\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]\leq 2D\sqrt{\frac{\sigma^{2}+(\delta+G)^{2}}{T}+\frac{\delta^{2}}{2}}\quad\text{(resp. $\frac{2D^{2}+\sigma^{2}+\delta^{2}+G^{2}}{\sqrt{T}}+(2D^{2}+1)\delta$ ).}$
@@
Let $w_{*}\in{\rm{argmin}}$$\ell$ and $T\geq 1$ be fixed. Develop the squares and use the definition of the algorithm to write
@@
Let $\bar{\bm{g}}_{t}$ be the projection of $\mathbb{E}\left[\bm{g}_{t}|\mathcal{F}_{t}\right]$ onto $\partial\ell(\bm{w}_{t})$, the subdifferential of $\ell$ at $\bm{w}_{t}$, which is a nonempty closed convex set. This means that from our assumption on the bias of $\bm{g}_{t}$ we will have
@@
Now we introduce a parameter $\varepsilon>0$, and we use Young’s inequality together with our assumption on the bias of $\bm{g}_{t}$ to write
$2\langle\bar{\bm{g}}_{t}-\mathbb{E}_{t}\left[\bm{g}_{t}\right],\bm{w}_{t}-\bm{w}_{*}\rangle\leq\frac{1}{\varepsilon}\|\bar{\bm{g}}_{t}-\mathbb{E}_{t}\left[\bm{g}_{t}\right]\|^{2}+\varepsilon\|\bm{w}_{t}-\bm{w}_{*}\|^{2}\leq\varepsilon^{-1}\delta^{2}+\varepsilon\|\bm{w}_{t}-\bm{w}_{*}\|^{2}.$
@@
$\mathbb{E}_{t}\left[\|\bm{w}_{t+1}-\bm{w}_{*}\|^{2}\right]\leq(1+\eta\varepsilon)\|\bm{w}_{t}-\bm{w}_{*}\|^{2}-2\eta(\ell(\bm{w}_{t})-\inf\ell)+\eta^{2}\sigma^{2}+\eta^{2}(\delta+G)^{2}+\eta\varepsilon^{-1}\delta^{2},$
@@
$\mathbb{E}\left[\ell(\bm{w}_{t})-\inf\ell\right]\leq\frac{1+\eta\varepsilon}{2\eta}\mathbb{E}\left[\|\bm{w}_{t}-\bm{w}_{*}\|^{2}\right]-\frac{1}{2\eta}\mathbb{E}\left[\|\bm{w}_{t+1}-\bm{w}_{*}\|^{2}\right]+\eta\frac{\sigma^{2}+(\delta+G)^{2}}{2}+\varepsilon^{-1}\frac{\delta^{2}}{2}.$
We are now ready to use a weighted telescopic sum argument. First, define $\theta:=\frac{1}{1+\eta\varepsilon}$ and $\rho_{t}:=\theta^{t}$. Second, multiply our inequality by $\rho_{t+1}$ and then sum for $t=0,\dots,T-1$. Observe that the term $(1+\eta\varepsilon)\rho_{t+1}$ is equal to $\rho_{t}$ due to our definition, which means that we have a telescoping sum where the terms $\tfrac{\rho_{t}}{2\eta}\mathbb{E}\left[\|\bm{w}_{t}-\bm{w}_{*}\|^{2}\right]$ will cancel each other. Third, divide by $\sum_{t=0}^{T-1}\rho_{t+1}$ so to obtain
$\frac{1}{\sum_{t=0}^{T-1}\rho_{t+1}}\sum_{t=0}^{T-1}\rho_{t+1}\mathbb{E}\left[\ell(\bm{w}_{t})-\inf\ell\right]\leq\frac{\rho_{0}}{2\eta\sum_{t=0}^{T-1}\rho_{t+1}}\mathbb{E}\left[\|\bm{w}_{0}-\bm{w}_{*}\|^{2}\right]+\eta\frac{\sigma^{2}+(\delta+G)^{2}}{2}+\varepsilon^{-1}\frac{\delta^{2}}{2}.$
@@
$\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]\leq\frac{\|\bm{w}_{0}-\bm{w}_{*}\|^{2}}{2\eta\sum_{t=0}^{T-1}\rho_{t+1}}+\eta\frac{\sigma^{2}+(\delta+G)^{2}}{2}+\varepsilon^{-1}\frac{\delta^{2}}{2}.$
We will now simplify this upper bound, so we can later make an appropriate choice of $\eta$ leading to the desired complexity bound. We will now focus on the geometric sum appearing in the denominator, and lower bound it. Start by using the definition of $\rho_{t}$ to write
@@
$\eta\frac{\theta}{1-\theta}=\frac{\eta}{1+\eta\varepsilon}\frac{1}{1-\frac{1}{1+\eta\varepsilon}}=\frac{\eta}{\eta\varepsilon}=\varepsilon^{-1}.$
On the other hand, we can guarantee that $(1-\theta^{T})\geq 1/2$ provided that $\varepsilon=\frac{2}{\eta T}$. To see this, first observe that $(1-\theta^{T})\geq 1/2$ is equivalent to $T\geq\frac{\ln(2)}{\ln(\theta^{-1})}$. Since we impose that $\eta\varepsilon=\frac{2}{T}$ then necessarily $\theta=\frac{T}{T+2}$ which means that $\theta^{-1}=1+\frac{2}{T}$. So what we need to verify is $T\geq\frac{\ln(2)}{\ln(1+\frac{2}{T})}$. Now, observe that $\ln(2)\leq 1$. Moreover, it is an exercise (see Lemma˜A.5 which is deferred to the end of this proof) to verify that for all $x\geq 0$, $\frac{1}{\ln(x+1)}\leq\frac{1}{2}+\frac{1}{x}$. So for our condition to hold it is enough to ask that $T\geq\frac{1}{2}+\frac{T}{2}$, which is equivalent to $T\geq 1$, which is true.
Combining all those bounds together with our new definition for $\varepsilon$, we now have
@@
Minimizing the right-hand side with respect to $\eta$ is equivalent to solve $\frac{a}{\eta T}=\eta b+c\eta T$, where
@@
We see that $\phi^{\prime\prime}(x)\geq 0$ for all $x\geq 0$, so $\phi$ is convex on $0,+\infty)$. So we can use the tangent inequality:
@@
A.4 Rates for SMGD for Convex Lipschitz Functions: Proof of Proposition˜5.4
See 5.4
@@
We want to apply Theorem˜A.4. Let us denote $\bm{g}_{t}=\bm{m}_{t}+\bm{e}_{t}$, so that the iterates of SMGD verify $\bm{w}_{t+1}=\bm{w}_{t}-\eta\bm{g}_{t}$. We are now going to show that $\bm{g}_{t}$ has uniformly bounded bias and variance. For this we will make use of Assumptions˜5.1 and 5.2. Remember that we assumed $\delta_{2}=\sigma_{2}=0$ without loss of generality (we can do this according to Lemma˜A.1), so we note $\delta$ and $\sigma$ instead of $\delta_{1}$ and $\sigma_{1}$. Also, Assumption˜5.2 on the errors together with the tower rule imply that $\mathbb{E}_{t}\left[\|\bm{e}_{t}\|^{2}\right]=\mathbb{E}_{t}\left[\mathbb{E}\left[\|\bm{e}_{t}\|^{2}\ |\ \bm{g}_{t}^{(1)},\dots,\bm{g}_{t}^{(n)}\right]\right]\leq\varepsilon^{2}$. Regarding the bias, we can write
@@
Using Theorem˜A.4, we obtain
@@
A.5 Rates for SMGD with Stable Noise: Proof of Corollary˜5.6
See 5.6
@@
First of all, remember that the $\ell_{1}$-median can be computed exactly, by computing component-wise a one-dimensional median. This allows us to have $\varepsilon=0$ in Assumption˜5.2. Further, due to Lemma˜A.3 we have $\delta_{1}=0$ in Assumption˜5.1.
Proof of part (i). Choose $\eta=\frac{1}{8L(1+2\sigma_{2}+2\delta_{2})\sqrt{T}}$, where $\delta_{2},\sigma_{2}$ are from Lemma˜A.3. Hence, we can apply Proposition˜5.3. As $\eta=\frac{C}{\sqrt{T}}$ with $C=\frac{1}{8L(1+2\sigma_{2}+2\delta_{2})}$, we get the bound $\frac{1}{C\sqrt{T}}+\frac{C}{\sqrt{T}}=\mathcal{O}(1/\sqrt{T})$.
Proof of part (ii). The result is a direct consequence of Lemma˜A.3 and Proposition˜5.4.
@@
Appendix B Approximate Computation of Gradient Estimators
B.1 Gradient Estimation Problem: General Considerations
@@
The function $\mathcal{D}:\mathbb{R}^{d}\to\mathbb{R}$ is convex, $\underset{\bm{z}}{\operatorname{argmin}}\;\mathcal{D}(\bm{z})=\{\bm{0}\}$ and problem (5) admits a solution.
Note that assuming $\mathcal{D}$ to be convex and finite implies that it is continuous, therefore measurable, which means that $\mathbb{E}\left[\mathcal{D}(\bm{m}-\bm{z})\right]$ is well defined. The assumption that $\underset{\bm{z}}{\operatorname{argmin}}\;\mathcal{D}(\bm{z})=\{\bm{0}\}$ guarantees that in the trivial case when the distribution is a Dirac $\delta_{\bm{m}_{0}}$, then the solution to (5) is exactly $\bm{m}_{0}$. The assumption that a solution to (5) exists is easily verified in practice, under the assumption that $\mathcal{D}$ is coercive:
@@
Let $\mathcal{D}:\mathbb{R}^{d}\to\mathbb{R}$ be convex. Assume that $\mathbb{E}\left[\mathcal{D}(\bm{z}-\mathbb{E}\left[\bm{z}\right])\right]<+\infty$. If $\mathcal{D}$ is coercive, i.e. $\lim\limits_{\|\bm{m}\|\to\infty}\mathcal{D}(\bm{m})=+\infty$, then (5) admits a solution.
@@
Let $\phi(\bm{m}):=\mathbb{E}\left[\mathcal{D}(\bm{m}-\bm{z})\right]$. Since we assume that $\phi(\mathbb{E}\left[\bm{z}\right])=\mathbb{E}\left[\mathcal{D}(\bm{z}-\mathbb{E}\left[\bm{z}\right])\right]<+\infty$ then $\phi$ is proper. Since $\mathcal{D}$ is convex then $\phi$ is convex, as an expectation of convex functions. Moreover $\mathcal{D}$ is convex and finite, which means that $\mathcal{D}$ is continuous on $\mathbb{R}^{d}$, see e.g. Corollary 8.39 in Bauschke & Combettes (2017). Therefore $\phi$ is an expectation of lower semicontinuous functions, which implies that $\phi$ is lower semicontinuous.
Since $\mathcal{D}$ is coercive, convex and continuous, we know that there exists $a\in(0,+\infty)$ and $b\in\mathbb{R}$ (see Proposition 14.16 in Bauschke & Combettes (2017)) such that $\mathcal{D}(\cdot)\geq a\|\cdot\|+b$. Therefore
@@
where in the inequalities we used the fact that $\|\cdot\|$ is $1$-Lipschitz, and that $a\mathbb{E}\left[\|\bm{z}-\mathbb{E}\left[\bm{z}\right]\|\right]\leq a\mathbb{E}\left[\mathcal{D}(\bm{z}-\mathbb{E}\left[\bm{z}\right])\right]-b<+\infty$. So $\phi$ is coercive on top of being proper convex and lower semicontinuous. We can then conclude that $\phi$ has a minimizer, with for instance Proposition 11.15 from Bauschke & Combettes (2017). ∎
B.2 Approximations Based on the Stochastic Proximal Point: General $\mathcal{D}$